### v2

# Fair Data Exchange at Near-Plaintext Efficiency

Abstract
The Fair Data Exchange (FDE) protocol (CCS ’24) provides atomic pay-per-file transfers with constant-size proofs, yet existing implementations remain unscalably slow ($\approx 1$ second per 4 KiB) and inflate ciphertexts by 10–50$\times$. We introduce an FDE implementation that achieves near-plaintext speeds and sizes, making fair exchange practical even for gigabyte-scale files.
Our approach leverages two key insights. First, we observe that a KZG commitment to polynomial evaluations implicitly (and without modification) also binds to the Reed–Solomon (RS) codeword of its coefficients, enabling sound and efficient randomized verification. Second, while heavyweight encryption schemes such as exponential ElGamal enable compact proofs linking ciphertexts to the commitment, they are unnecessary for direct data recovery.
Exploiting these insights, we apply a lightweight hash-derived mask to the entire RS-extended codeword, and perform ElGamal encryption only on a pseudorandom $\Theta(\lambda)$ subset of symbols, where $\lambda$ is the security parameter (e.g., 128). Data recovery occurs by simply removing the lightweight masks, with ElGamal ciphertexts serving exclusively for verification proofs. A heavyweight (but constant-time) zk-SNARK ensures consistency between these two encryption layers at sampled positions, sharply reducing bandwidth overhead and computational cost.
In addition, we show how a constant-time (and precomputable) zk-SNARK linking a BLS12-381 secret key to a secp256k1 hash pre-image resolves Bitcoin’s elliptic-curve mismatch, enabling fully off-chain execution via the Lightning Network. This can reduce transaction fees from roughly \$10 to under \$0.01 and shortens transaction latency from tens of seconds on Ethereum down to about a second or less.
1 Introduction
Digital commerce increasingly revolves around the one-shot sale of large digital assets—scientific data sets, proprietary machine-learning weights, high-resolution media, even genomic archives—between parties with no prior trust relationship. Ensuring that the buyer actually receives the promised file while the seller simultaneously receives payment is therefore a foundation stone of the data-driven economy. Traditional escrow or licensing services resolve this tension by inserting costly legal or institutional intermediaries; blockchain-based atomic fair-exchange protocols aim to provide the same fairness guarantee with only a smart contract as referee. Achieving that goal without inflating on-chain fees or forcing the parties through many interactive rounds, however, remains challenging once file sizes grow into the multi-megabyte range.
The prominent blockchain-based fair-exchange protocols—FairSwap [1], FileBounty [3], and FairDownload [2]—all follow a dispute-driven model, resolving misbehavior through an explicit on-chain arbitration phase. FairSwap requires both parties to remain responsive throughout a complaint period, during which buyers must provide Merkle proofs of misbehavior. FileBounty relaxes continuous responsiveness by allowing zk-SNARK-based disputes at any point, limiting potential loss to individual file chunks but implicitly assuming partial data has proportional utility. FairDownload resolves disputes via off-chain exchange of signed data chunks, with disputes settled on-chain using $O(\log k)$-sized Merkle inclusion proofs.
The FDE protocol [5] entirely eliminates dispute timers and achieves constant-sized on-chain communication. In FDE, the seller publishes a single KZG polynomial commitment [4] and provides an upfront, constant-sized Verifiable Encryption of Committed Knowledge (VECK) proof that every ciphertext sent to the buyer encrypts the correct polynomial evaluation. The protocol requires only three on-chain messages—commitment $\rightarrow$ payment $\rightarrow$ decryption key. The buyer can verify the VECK proof immediately, pay, and decrypt the file without waiting or worrying about subsequent disputes.
The core barrier limiting practical deployment of FDE is its substantial computational overhead. The VECK proofs introduced by Taş et al. [5] incur high computational costs for both prover and verifier. Even for small files (e.g., 128 KiB), generating the VECK proof alone requires over 40 seconds, while verification takes approximately 34 seconds on commodity hardware. [^1] Figure 6 in [5]; measurements for the ElGamal-based VECK protocol. Extrapolating these runtimes, exchanging just a 128 MiB file would demand roughly 11 hours of proof generation and over 9 hours of verification, rendering large-scale deployments impractical for common use-cases such as AI model distribution or genomic data transfers.
Communication overhead further compounds this challenge. The ElGamal- and Paillier-based implementations in [5] inflate ciphertext sizes by 10–50$\times$ relative to the original data, significantly increasing bandwidth usage and limiting practicality for large-file transfers.
Our contributions. In this work, we introduce significant improvements to FDE, reducing computational and communication costs dramatically, making it practical for gigabyte-scale files while maintaining its original strong fairness and atomicity guarantees. Our enhancements are driven by two key insights:
First, we observe that when a file is represented as evaluations of a polynomial $\phi$, the original KZG polynomial commitment implicitly—and without modification—also binds to any Reed–Solomon (RS) codeword obtained by extending the polynomial’s coefficients with a constant fraction of redundancy. Crucially, this insight enables us to apply our coding-based verification directly to files for which a commitment already exists, instead of first encoding the file and then applying the commitment. The alternative approach—committing directly to the coded file—would, especially in constrained blockchain environments (e.g., Ethereum blobs limited to 128 KiB), significantly reduce the effective blob space available for useful data and obligate validators to perform additional computations to verify correct encoding.
Second, we recognize that heavyweight encryption schemes such as exponential ElGamal are necessary only for generating a small set of ciphertexts to verify consistency against the commitment, but are unnecessary—and computationally wasteful—for bulk data recovery. Once the buyer obtains the secret key, lightweight masking with a hash-derived stream cipher suffices for efficient file recovery. Crucially, eliminating ElGamal encryption from the recovery step removes the need to partition data into small chunks for brute-force decoding (as done previously in [5]), avoiding substantial ciphertext expansion and large memory overhead required by precomputed lookup tables.
Technical approach. Building upon these insights, our approach encodes the file with a constant-rate Reed–Solomon redundancy, masks the entire extended codeword with a fast, hash-derived cipher, and applies ElGamal encryption only to a pseudorandom $\Theta(\lambda)$ subset of symbols, where $\lambda$ is the 128-bit security parameter. A heavyweight (but constant-time) zk-SNARK establishes cryptographic consistency between the lightweight masked data and the sampled ElGamal ciphertexts. After payment, the buyer removes the lightweight mask, verifies correctness on the sampled subset, and employs RS decoding if necessary to correct any residual corruptions, ensuring negligible failure probability. Because public-key ciphertexts and associated proofs appear only at $\Theta(\lambda)$ positions, the overall computational complexity collapses from $O(n)$ to $O(\lambda)$, and bandwidth overhead shrinks significantly.
Bulk encryption and decryption of the file now leverage the lightweight, hash-based masking method, enabling extremely fast symmetric-key operations. By eliminating reliance on heavyweight public-key operations for most of the data, we substantially reduce both runtime and memory requirements. Taken together, these improvements allow our protocol to achieve performance and bandwidth overhead approaching that of plaintext data transfers, even for multi-gigabyte files.
Bitcoin and Lightning compatibility. Additionally, we identify and resolve a critical elliptic-curve incompatibility between Bitcoin’s secp256k1-based adaptor signatures and the BLS12-381-based VECK proofs required by the original FDE design. We introduce a compact, constant-time zk-SNARK bridging these curves, enabling the fair exchange to execute entirely off-chain over the Lightning Network. This significantly reduces transaction fees from roughly \$10 per exchange on Ethereum to under \$0.01, while cutting latency from tens of seconds (due to multiple sequential Ethereum transactions) to about a second or less, substantially enhancing practicality for large-scale deployments.
2 Background
2.1 Notation and Definitions
Let $\lambda\in\mathbb{N}$ denote the security parameter. A non-negative function $\sigma(\lambda)$ is said to be negligible if, for every polynomial $p(\lambda)$, there exists a sufficiently large $\lambda_{0}$ such that for all $\lambda\geq\lambda_{0}$:
$\sigma(\lambda)\leq\frac{1}{p(\lambda)}.$
For a random variable $x$, we write $x\leftarrow_{R}X$ to indicate that $x$ is drawn uniformly at random from the set $X$.
Throughout this paper, we denote by $\mathbb{F}_{p}$ the finite field of prime order $p$. We let $\mathbb{F}_{p}[X]$ represent the set of all univariate polynomials with coefficients in $\mathbb{F}_{p}$. Specifically, we denote by $\mathbb{F}_{p}^{k}[X]$ the set of polynomials of degree exactly $k$, and by $\mathbb{F}_{p}^{\leq k}[X]$ the set of polynomials of degree at most $k$.
We work with elliptic-curve groups $G_{1}$, $G_{2}$, and a target group $G_{T}$, each of prime order $p$. We utilize a cryptographic bilinear pairing map $e:G_{1}\times G_{2}\rightarrow G_{T}$, satisfying standard properties of bilinearity, non-degeneracy, and efficient computability. Our constructions specifically utilize the pairing-friendly elliptic curve BLS12-381, which provides approximately $128$ bits of security. The prime order $p$ in the BLS12-381 curve is roughly $255$ bits in length.
For a set $S$ and a function $\phi(X)\in\mathbb{F}_{p}[X]$, $\phi_{S}(X)$ denotes the minimal-degree polynomial in $\mathbb{F}_{p}[X]$ satisfying $\phi_{S}(i)=\phi(i)$ for all $i\in S$. Additionally, define
$V_{S}(X)\coloneqq\prod_{i\in S}(X-i).$
2.2 KZG Polynomial Commitments
The KZG polynomial commitment scheme [4] provides a cryptographic method for succinctly committing to polynomials and efficiently generating evaluation proofs. It leverages pairing-based cryptography to achieve constant-size commitments and evaluation proofs, independently of the polynomial degree.
Formally, the KZG polynomial commitment scheme consists of the following algorithms:
  • $\mathsf{Setup}(1^{\lambda},n)\rightarrow\mathsf{crs}$: generates $\mathcal{G}$: elliptic-curve groups $G_{1},G_{2},G_{T}$ of prime order $p\geq 2^{2\lambda}$, with generators $g_{1}\in G_{1}$, $g_{2}\in G_{2}$, $g_{T}\in G_{T}$, and a bilinear pairing map $e:G_{1}\times G_{2}\rightarrow G_{T}$. Samples a uniformly random secret $\tau\leftarrow_{R}\mathbb{F}_{p}$ and publish
    $\mathsf{crs}=\bigl{(}\mathcal{G},\{g_{1}^{\tau^{i}}\}_{i=1}^{n},\{g_{2}^{\tau^{i}}\}_{i=1}^{n}\bigr{)}.$
  • $\mathsf{Commit(\mathsf{crs},\phi)}\rightarrow C$: Given the $\mathsf{crs}$ and the coefficients of $\phi(X)\in\mathbb{F}_{p}[X]$, output the commitment
    $C=g_{1}^{\phi(\tau)}.$
  • $\mathsf{Open}(\mathsf{crs},\phi,i)\rightarrow\pi$: Given public parameters $\mathsf{crs}$, polynomial $\phi(X)\in\mathbb{F}_{p}[X]$, and an evaluation point $i$, the prover outputs the opening proof:
    $\pi=g_{1}^{\frac{\phi(\tau)-\phi(i)}{\tau-i}}\in G_{1}.$
  • $\mathsf{Verify}(\mathsf{crs},C_{\phi},i,\phi(i),\pi)\rightarrow 0/1$: To verify an opening, the verifier checks the pairing equation:
    $e(C_{\phi}/g_{1}^{\phi(i)},g_{2})\stackrel{{\scriptstyle?}}{{=}}e(\pi_{i},g_{2}^{\tau-i}),$
    and outputs 1 if it holds, otherwise 0.
  • $\mathsf{batchOpen}(\mathsf{crs},\phi,S)\rightarrow\pi$: Given public parameters $\mathsf{crs}$, polynomial $\phi(X)$, and multiple distinct evaluation points $S$, the prover computes the evaluations $\{\phi(i)\}_{i\in S}$ and produces a single aggregated proof:
    $\pi=g_{1}^{q(\tau)},$
    where
    $q(X)=\frac{\phi(X)-\phi_{S}(X)}{V_{S}(X)}.$
  • $\mathsf{batchVerify}(\mathsf{crs},C_{\phi},S,\{\phi(i)\}_{i\in S},\pi)\rightarrow 0/1$: The verifier efficiently checks the aggregated proof $\pi$ by verifying the following pairing equation:
    $e\left(C_{\phi}/g_{1}^{\phi_{S}(\tau)},\,g_{2}\right)\stackrel{{\scriptstyle?}}{{=}}e\left(\pi,\,g_{2}^{V_{S}(\tau)}\right).$
    The verifier outputs 1 if this equation holds and 0 otherwise.
2.3 FDE Protocol
FDE is a blockchain-based protocol which enables a client and a storage server to atomically exchange data for payment. Atomicity guarantees that the server receives payment if and only if the client obtains the promised data. The FDE protocol leverages the KZG polynomial commitment scheme, chosen due to its constant-size commitments and efficient batchable opening proofs, making it particularly suitable for scenarios where clients may retrieve subsets of data. Additionally, KZG commitments are already widely adopted in blockchain ecosystems, notably in Ethereum’s Danksharding for data availability, making FDE naturally compatible with existing infrastructure.
In the FDE protocol, data is represented as evaluations of a polynomial $\phi(\cdot)$ of degree $\ell\leq n$, that is as $\{\phi(i)\}_{i=0}^{\ell}$. The server first publishes a public verification key $\mathsf{vk}$ to a blockchain smart contract, alongside specific transaction details such as the agreed price and the client’s blockchain address (step 1). Off-chain, the server then sends the encrypted evaluations $\{\mathsf{ct}_{i}\}_{i=0}^{\ell}$ of each data point $\{\phi(i)\}_{i=0}^{\ell}$ to the client. These ciphertexts are accompanied by a cryptographic proof showing that each ciphertext $\mathsf{ct}_{i}$ correctly encrypts the corresponding polynomial evaluation $\phi(i)$ committed to by a KZG polynomial commitment $C_{\phi}$, under a secret decryption key $\mathsf{sk}$ that matches the previously submitted verification key $\mathsf{vk}$ (step 2).
Upon receiving and verifying these ciphertexts and associated proofs, the client locks the agreed-upon funds in the blockchain smart contract (step 3). The server subsequently can claim these funds only by revealing the correct decryption key $\mathsf{sk}$ that corresponds to the public verification key $\mathsf{vk}$ (step 4). After the server publishes the decryption key, the client retrieves it from the blockchain (step 5), allowing immediate decryption of the ciphertexts and recovery of the original committed data (step 6). If the server fails to reveal the secret key within a specified timeout, the client recovers the funds locked in the contract, thus preserving atomicity.
FDE satisfies three critical properties: correctness (honest parties always succeed), client-fairness (the server obtains payment only if the client receives the data), and server-fairness (the client learns nothing about the data unless the server is paid).
2.4 Verifiable Encryption under Committed Key (VECK)
At the heart of the FDE protocol lies a novel cryptographic primitive called VECK. At a high level, VECK enables a prover (in our context, the server) to demonstrate that a set of ciphertexts indeed encrypts evaluations of a polynomial at specific points, using an encryption key consistent with a publicly committed verification key. Concretely, VECK allows efficient verification of ciphertext correctness against a polynomial commitment without revealing the underlying plaintext evaluations or the encryption key itself. This construction ensures that the verifier (client) learns no additional information beyond the correctness of ciphertexts until the prover reveals the corresponding decryption key.
Formal description. Let $(\mathsf{Setup},\mathsf{Commit})$ be a non-interactive binding commitment scheme, where $\mathsf{Setup}(1^{\lambda},n)\rightarrow\mathsf{crs}$ generates a public common reference string, and $\mathsf{Commit}(\mathsf{crs},w\in W)\rightarrow C_{w}$ generates a commitment to $w$. A non-interactive VECK scheme for a class of functions =$\mathcal{F}=\{F:W\rightarrow V\}$ is defined as follows [5]:
  • $\mathsf{VECK.Gen(\mathsf{crs})}\rightarrow\mathsf{pp}$: A PPT algorithm that, given the $\mathsf{crs}$, outputs parameters $\mathsf{pp}$ and defines relevant spaces. The parameters $\mathsf{pp}$ are implicitly provided to subsequent algorithms.
  • $\mathsf{VECK.Enc(F,C_{w},w)}\rightarrow(\mathsf{vk},\mathsf{sk},\mathsf{ct},\pi)$: A PPT algorithm run by the server that takes $\mathsf{(F,C_{w},w)}$ and outputs a verification key $\mathsf{vk}$, a decryption key $\mathsf{sk}$ and the encryptions $\mathsf{ct}$ of $F(w)$, as well as a proof $\pi$.
  • $\mathsf{VECK.Ver_{ct}(F,C_{w},vk,ct,\pi)}\rightarrow\{0,1\}$: A deterministic polynomial-time algorithm executed by the client that returns either accept or reject.
  • $\mathsf{VECK.Ver_{key}(vk,sk)}\rightarrow\{0,1\}$: A deterministic polynomial-time algorithm executed by the blockchain or a trusted third party to verify the validity of the secret key.
  • $\mathsf{VECK.Dec(\mathsf{sk},\mathsf{ct})}\rightarrow v/\bot$: A deterministic polynomial-time algorithm executed by the client that outputs a value in $V$ or a failure symbol $\bot$.
A secure VECK scheme must satisfy the following key properties:
  • Correctness (informal): Honestly generated ciphertexts and proofs always verify correctly, and decryption recovers the original polynomial evaluations.
  • Soundness (informal): No computationally bounded adversary can generate ciphertexts and corresponding proofs that pass verification yet encrypt values different from the committed polynomial evaluations, unless cryptographic assumptions are broken.
  • Zero-Knowledge (informal): The proof and ciphertexts do not reveal any additional information about the underlying polynomial evaluations or secret encryption key beyond correctness until the decryption key is explicitly revealed.
2.5 Reed–Solomon Codes
Let $(a_{0},\dots,a_{\ell})\in\mathbb{F}_{p}^{\ell+1}$ be a message encoded by the polynomial
$\phi(X)=\sum_{i=0}^{\ell}a_{i}X^{\,i}.$
Given $n\geq\ell+1$ pairwise-distinct evaluation points $\alpha_{1},\dots,\alpha_{n}\in\mathbb{F}_{p}$, the corresponding Reed–Solomon (RS) codeword is
$\mathbf{c}=(\phi(\alpha_{1}),\dots,\phi(\alpha_{n}))\in\mathbb{F}_{p}^{\,n}.$
This yields an $(n,\ell+1)$ linear code with rate $R=(\ell+1)/n$ and minimum distance $d_{\min}=n-\ell$, capable of correcting up to $t=\lfloor(d_{\min}-1)/2\rfloor$ symbol errors.
Fast deterministic detection via the parity check. Let $H\in\mathbb{F}_{p}^{(n-k)\times n}$ be the parity-check matrix of the code ($k=\ell+1$). For a received word $\mathbf{c}=(c_{1},\dots,c_{n})$, compute the syndrome vector
$\mathbf{s}=H\,\mathbf{c}^{\mathsf{T}}\in\mathbb{F}_{p}^{\,n-k}.$
This requires only a single $O(n)$ streaming pass. If $\mathbf{s}=0$, the word is already a valid codeword and full decoding can be skipped; otherwise the decoder is invoked. This detect-then-correct strategy makes the common error-free case inexpensive.
Optional probabilistic optimization (single accumulator). We apply the standard random linear-combination (Freivalds) test to compress the syndrome check to a single scalar. Fix a random non-zero vector $v\in\mathbb{F}_{p}^{\,n-k}$ once and pre-compute
$w=v^{\mathsf{T}}H\in\mathbb{F}_{p}^{\,n}.$
In one streaming pass, compute
$T=\sum_{j=1}^{n}w_{j}\,c_{j}.$
If $T=0$ we accept immediately; otherwise full decoding is performed. A non-codeword passes this test with probability exactly $1/p$ (e.g., $2^{-255}$ in the 255-bit BLS12-381 field), while the online cost drops to one field multiplication and one addition per symbol (about 32 bytes of data).
3 Proposed FDE Implementations
Our primary goal is to make FDE efficient and practical for large-scale files. To achieve this, we propose two alternative implementations of the FDE protocol: the first leverages RS codes to substantially reduce proof complexity and runtime, while the second further enhances computational efficiency and minimizes communication overhead through optimized masking techniques combined with constant-time zk-SNARK proofs.
These implementations rely on two novel ElGamal-based VECK protocols, denoted as $\mathsf{VECK^{+}_{EL}}$ and $\mathsf{VECK^{\star}_{EL}}$. Before detailing these implementations, we briefly introduce notation and definitions common to both implementations.
3.1 Common Notation and Definitions
As in [5], we use the KZG polynomial commitment scheme, and define the function $F$ in VECK as the polynomial evaluation at a given set of points.
Our first proposed construction, $\mathsf{VECK^{+}_{EL}}$, uses ElGamal-based encryption. We choose ElGamal primarily because it achieves significantly lower bandwidth compared to the Paillier-based VECK from [5], despite its relatively higher proof-time complexity. This trade-off is justified in $\mathsf{VECK^{+}_{EL}}$, as our design already reduces proof generation time to essentially constant complexity, making the bandwidth savings more impactful.
Our second construction, $\mathsf{VECK^{\star}_{EL}}$, also employs ElGamal-based encryption, primarily to maintain consistency between the two implementations. However, unlike in $\mathsf{VECK^{+}_{EL}}$, the bandwidth advantage of ElGamal over Paillier is less critical, as $\mathsf{VECK^{\star}_{EL}}$ already drastically reduces communication overhead through optimized masking and zk-SNARK integration. Thus, in the context of $\mathsf{VECK^{\star}_{EL}}$, one might prefer Paillier encryption to leverage its slightly better computational efficiency, as reported in [5].
Let $\mathsf{VECK_{El}}$ denote the Elgamal-based VECK protocol proposed in [5]. Since our VECK implementations have similarities with $\mathsf{VECK_{El}}$, instead of constructing them from scratch, we strive to call $\mathsf{VECK_{El}}$ operation whenever possible to reduce repetition of basic operations already covered in the construction of $\mathsf{VECK_{El}}$. Towards this end, we split the operations $\mathsf{VECK_{El}.Enc}\rightarrow\mathsf{(vk,sk,ct,\pi)}$ into two sub-operations:
$\mathsf{VECK_{El}.Enc_{1}}\rightarrow\mathsf{(vk,sk,ct)}$
and
$\mathsf{VECK_{El}.Enc_{2}}\rightarrow(\pi,\mathsf{ct}_{-}).$
The first sub-operation returns all outputs except the proof $\pi$, while the second sub-operation returns only the proof along with $\mathsf{ct}_{-}$, an encryption at point $-1$ employed in $\mathsf{VECK_{El}.Enc}$ to assist with ciphertext verification.
Furthermore, we stipulate that if $\mathsf{VECK_{El}.Dec}$ fails to recover a particular value (i.e., the plaintext lies outside the brute-force search range), it returns $\bot$ for that specific position only, which we interpret as an erasure.
We employ Reed–Solomon detection and decoding as black-box subroutines defined as follows:
$\mathsf{RS.Det}(S,\{\tilde{\phi}_{i}\}_{i\in S})\;\longrightarrow\;\{0,1\}\quad\text{and}\quad\mathsf{RS.Dec}(S^{\prime},\,S,\,\{\tilde{\phi}_{i}\}_{i\in S})\;\longrightarrow\;\{\phi(i)\}_{i\in S^{\prime}}\cup\{\bot\},$
where $S\subseteq\mathbb{F}_{p}$ is an ordered set of evaluation points, $S^{\prime}$ is the target set (typically a subset of $S$), and $\{\tilde{\phi}_{i}\}_{i\in S}$ denotes the ordered sequence of received symbols, which may potentially be corrupted.
Specifically, $\mathsf{RS.Det}$ checks whether the received sequence is a valid RS codeword, returning 00 if it is valid and $1$ otherwise. The decoding algorithm $\mathsf{RS.Dec}$ takes the (potentially corrupted) received sequence and attempts to reconstruct the polynomial evaluations at the points in the subset $S^{\prime}$. If decoding succeeds, it outputs the corrected evaluations $\{\phi(i)\}_{i\in S^{\prime}}$; otherwise, it returns $\bot$.
3.2 FDE via $\mathsf{VECK^{+}_{EL}}$
Our first implementation leverages the observation that when a file is represented by evaluations of a polynomial $\phi$, its original KZG polynomial commitment implicitly (and without modification) also binds to any RS codeword formed directly from the polynomial’s coefficients. This insight enables coding-based verification directly on existing commitments, without additional on-chain commitments or overhead.
Case $S=[\ell]$. This corresponds to the scenario in which the client requests the entire data set. We handle the case $S\subset[\ell]$ separately, as requesting the full data set ($S=[\ell]$) allows fewer steps in both $\mathsf{VECKplus_{El}.Enc}$ and $\mathsf{\mathsf{VECK^{+}_{EL}}.Ver_{ct}}$.
$\mathsf{\mathsf{VECK^{+}_{EL}}}$:
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.GEN(crs)}\rightarrow\mathsf{pp}$: On input $\mathsf{crs}=\bigl{(}\mathcal{G},\{g_{1}^{\tau_{i}}\}_{i=1}^{n},\{g_{2}^{\tau_{i}}\}_{i=1}^{n}\bigr{)}$, generate random group elements with unknown discrete logarithms $h\leftarrow_{R}G_{1}$ and $h_{i}\leftarrow_{R}G_{1}$ for $i\in[m]\cup\{-1\}$, where $m=\lceil\beta\cdot\ell\rceil$.
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.Enc(F_{[\ell]},C_{\phi},\phi(X)})\rightarrow\mathsf{(vk,sk,ct,\pi)}$.
    1. 1.
      Compute $\mathsf{(vk,sk,ct)}\coloneqq\mathsf{VECK_{El}.Enc_{1}(F_{[m]},C_{\phi},\phi(X))}$
    2. 2.
      Generate a Fiat-Shamir pseudo-random challenge subset $S_{R}\subset[m]$ with $|S_{R}|=\min(\ell+1,\lceil\frac{\lambda}{\beta-1}\rceil)$.
    3. 3.
      Compute $(\pi_{R},\mathsf{ct}_{-})\coloneqq\mathsf{VECK_{El}.Enc_{2}(F_{S_{R}},C_{\phi},\phi(X))}.$
    4. 4.
      Output $\mathsf{(vk,sk,ct,\pi=\mathsf{(\pi_{R},ct_{-})}})$
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.Ver_{ct}(F_{[\ell]},C_{\phi},vk,ct,\pi)}\rightarrow 0/1$
    1. 1.
      Parse $\pi$ as $(\pi_{R},\mathsf{ct}_{-})$.
    2. 2.
      Output $\mathsf{VECK_{El}.Ver_{ct}(F_{S_{R}},C_{\alpha},vk,\{ct_{i}\}_{i\in S_{R}}\cup ct_{-},\pi_{R})}$.
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.Ver_{key}(vk,sk)}\rightarrow 0/1:$ For $\mathsf{sk}=s\in\mathbb{F}_{p}$, return 1 iff $\mathsf{vk}=h^{s}$.
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.Dec(F_{[m]},sk,ct)}\rightarrow\{\phi(i)\}_{i\in[\ell]}$.
    1. 1.
      Compute $\{\tilde{\phi}(i)\}_{i\in[m]}\coloneqq\mathsf{VECK.Dec(F_{[m]},sk,ct)}$.
    2. 2.
      If $\mathsf{RS.Det}([m],\{\tilde{\phi}(i)\}_{i\in[m]})=0$ (no error detected), then output $\{\tilde{\phi}(i)\}_{i\in[\ell]}$. Otherwise, run the decoding algorithm and output
      $\mathsf{RS.Dec}\left([\ell],[m],\{\tilde{\phi}(i)\}_{i\in[m]}\right).$
Theorem 3.1
For the case $S=[\ell]$, the generation and verification times of $\mathsf{VECK^{+}_{EL}}$ are $\mathcal{O}(\lambda)$.
Proof
Each step (specifically, steps 2 and 3) in the proof generation runs in $\mathcal{O}(\lambda)$ time: generating the random subset $S_{R}$ requires $\mathcal{O}(R)=\mathcal{O}(\lambda)$ operations, and $\mathsf{VECK_{El}.Enc_{2}}$ (step 3) has complexity linear in $|S_{R}|$, which is also $\mathcal{O}(\lambda)$. Similarly, the time complexity of $\mathsf{VECK_{El}.Ver_{ct}}$ is linear in $|S_{R}|\in\mathcal{O}(\lambda)$.
Case $S\subset[\ell]$. This corresponds to the scenario in which the client does not request the entire data set.
$\mathsf{\mathsf{VECK^{+}_{EL}}}$:
  • $\mathsf{VECK_{El}^{+}.GEN(crs)}\rightarrow\mathsf{pp}$: On input $\mathsf{crs}=\bigl{(}\mathcal{G},\{g_{1}^{\tau_{i}}\}_{i=1}^{n},\{g_{2}^{\tau_{i}}\}_{i=1}^{n}\bigr{)}$, generate random group elements with unknown discrete logarithms $h\leftarrow_{R}G_{1}$ and $h_{i}\leftarrow_{R}G_{1}$ for $i\in[m]\cup\{-1\}$, where $m=\lceil\beta\cdot|S|\rceil$.
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.Enc(F_{S},C_{\phi},\phi(X)})\rightarrow\mathsf{(vk,sk,ct,\pi)}$.
    1. 1.
      Sample $t\leftarrow_{R}\mathbb{F}_{p}$.
    2. 2.
      Set $\phi^{\prime}_{S}(X)\coloneqq t\,V_{S}(X)+\phi_{S}(X)$.
    3. 3.
      Compute $C_{S}\coloneqq\mathsf{Commit}(\mathsf{crs},\,\phi^{\prime}_{S}(X)).$
    4. 4.
      Compute $\mathsf{(vk,sk,ct)}\coloneqq\mathsf{VECK_{EL}.Enc_{1}(F_{[m]},C_{S},\phi^{\prime}_{S}(X))}$
    5. 5.
      Compute $\pi_{S}\coloneqq\mathsf{batchOpen}\bigl{(}\mathsf{crs},\,\phi(X)-\phi^{\prime}_{S}(X),S\bigr{)}.$
    6. 6.
      Generate a Fiat-Shamir pseudo-random challenge subset $|S_{R}|=\min(|S|+1,\lceil\frac{\lambda}{\beta-1}\rceil)$.
    7. 7.
      Compute $(\pi_{R},\mathsf{ct}_{-})\coloneqq\mathsf{VECK_{El}.Enc_{2}(F_{S_{R}},C_{S},\phi^{\prime}_{S}(X))}.$
    8. 8.
      Output $\mathsf{\bigl{(}vk,sk,ct,\pi=(\pi_{S},\pi_{R},ct_{-})\bigr{)}}$
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.Ver_{ct}(F_{S},C_{\phi},vk,ct,\pi)}\rightarrow 0/1$
    1. 1.
      Parse $\pi$ as $(\pi_{S},\pi_{R},\mathsf{ct}_{-})$.
    2. 2.
      Compute $b_{1}\coloneqq\mathsf{batchVefiry(crs,C_{\phi}/C_{S},S,\mathbf{0},\pi_{S})}$.
    3. 3.
      Compute $b_{2}=\mathsf{VECK_{El}.Ver_{ct}(F_{S_{R}},C_{S},vk,\{ct_{i}\}_{i\in S}\cup ct_{-},\pi_{R})}$.
    4. 4.
      Output $b_{1}\land b_{2}$.
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.Ver_{key}(vk,sk)}\rightarrow 0/1:$ For $\mathsf{sk}=s\in\mathbb{F}_{p}$, return 1 iff $\mathsf{vk}=h^{s}$.
  • $\mathsf{\mathsf{VECK^{+}_{EL}}.Dec(F_{[m]},sk,ct)}\rightarrow\{\phi(i)\}_{i\in S}$.
    1. 1.
      Compute $\{\tilde{\phi}(i)\}_{i\in[m]}\coloneqq\mathsf{VECK.Dec(F_{[m]},sk,ct)}$.
    2. 2.
      If $\mathsf{RS.Det}([m],\{\tilde{\phi}(i)\}_{i\in[m]})=0$ (no error detected), then output $\{\tilde{\phi}(i)\}_{i\in[\ell]}$. Otherwise, run the decoding algorithm and output
      $\mathsf{RS.Dec}\left(S,[m],\{\tilde{\phi}(i)\}_{i\in[m]}\right).$
Theorem 3.2
Let $S\subset[\ell]$.
  1. (i)
    If the verifier has the pre-computed element $g_{2}^{\,V_{S}(\tau)}$, the running time of $\mathsf{VECK^{+}_{El}.Ver_{ct}}$ is $\mathcal{O}(\lambda)$.
  2. (ii)
    Otherwise, the prover can supply a KZG opening proof for $V_{S}(X)$ at a Fiat–Shamir-chosen point $\kappa$, in which case verification costs $\mathcal{O}(|S|+\lambda)$.
Proof
When $g_{2}^{\,V_{S}(\tau)}$ is available, Step 2 requires only a constant number of group operations, i.e. $\mathcal{O}(1)$. Step 3 performs $\lambda$ pairing/exponentiation checks, so the total cost is $\mathcal{O}(\lambda)$.
If the verifier lacks $g_{2}^{\,V_{S}(\tau)}$, the prover can provide a KZG opening $\pi$ of $V_{S}(X)$ at $\kappa$. Verifying $\pi$ costs $\mathcal{O}(1)$ pairings, but the verifier must evaluate $V_{S}(\kappa)$ locally, which takes $\mathcal{O}(|S|)$ field operations. Together with the $\mathcal{O}(\lambda)$ checks, this yields $\mathcal{O}(|S|+\lambda)$ overall.
Theorem 3.3
In the random-oracle and algebraic-group models, the protocol $\mathsf{\mathsf{VECK^{+}_{EL}}}$ satisfies correctness, soundness, and computational zero-knowledge; hence it is a secure $\mathsf{VECK}$.
3.3 FDE via $\mathsf{VECK^{\star}_{EL}}$
Building on the RS-based approach, our second implementation improves efficiency by using a two-layer encryption strategy. Instead of applying expensive exponential ElGamal encryption directly to all symbols, we encrypt only a small pseudorandom subset. The entire extended RS-encoded file is protected with a lightweight masking method derived from a fast, hash-based stream cipher. A heavyweight (but constant-time) zk-SNARK then succinctly proves consistency between the masked data and the sampled, publicly encrypted symbols.
This second protocol significantly reduces computational complexity, communication overhead, and decoding latency. Since most symbols rely on symmetric-key masking rather than public-key encryption, decryption and verification approach near-plaintext efficiency, making FDE practical even for large, gigabyte-scale data transfers.
In the protocol, $H:\{0,1\}^{\ast}\!\to\!\mathbb{F}_{p}$ denotes a cryptographic hash function modeled as a random oracle; all additions involving $H$ are taken modulo $p$. We note that $\mathsf{VECK^{\star}_{EL}}$ inherently requires a much shorter common reference string ($\mathsf{crs}$), since it needs parameters only for the sampled subset $S_{R}$, whose size is typically much smaller than $n$. However, for consistency and clarity of presentation, we retain the same $\mathsf{crs}$ length across all implementations
Case $S=[\ell]$. This corresponds to the scenario where the client requests the entire data set. The alternative scenario, in which $S\subset[\ell]$, can be handled similarly to the procedure outlined earlier for $\mathsf{VECK^{+}_{EL}}$.
$\mathsf{\mathsf{VECK^{\star}_{EL}}}$:
  • $\mathsf{\mathsf{VECK^{\star}_{EL}}.GEN(crs)}\rightarrow\mathsf{pp}$: On input $\mathsf{crs}=\bigl{(}\mathcal{G},\{g_{1}^{\tau_{i}}\}_{i=1}^{n},\{g_{2}^{\tau_{i}}\}_{i=1}^{n}\bigr{)}$, generate random group elements with unknown discrete logarithms $h\leftarrow_{R}G_{1}$ and $h_{i}\leftarrow_{R}G_{1}$ for $i\in[m]\cup\{-1\}$, where $m=\lceil\beta\cdot\ell\rceil$.
  • $\mathsf{\mathsf{VECK^{\star}_{EL}}.Enc(F_{[\ell]},C_{\phi},\phi(X)})\rightarrow\mathsf{(vk,sk,ct,\pi)}$.
    1. 1.
      Sample $s\leftarrow_{R}\mathbb{F}_{p}$, set $\mathsf{sk}:=s$, $\mathsf{vk}:=h^{s}$.
    2. 2.
      Compute $\mathsf{ct}=\{\mathsf{ct}_{i}\coloneqq\phi(i)+H(\mathsf{sk},i)\}_{i\in[m]}$.
    3. 3.
      Generate the Fiat–Shamir challenge
      $S_{R}\;=\;\mathsf{DeriveSubset}\!\bigl{(}H\bigl{(}C_{\phi},\,\mathsf{vk},\,\mathsf{ct}\bigr{)},m,\;|S_{R}|\bigr{)},$
      where $|S_{R}|=\min\!\bigl{(}\ell+1,\lceil\lambda/(\beta-1)\rceil\bigr{)}$.
    4. 4.
[…]
    5. 5.
      Compute a zk-SNARK proof $\pi_{Z}$ for the relation $\mathcal{R}$:
      =$\mathcal{R}\;=\;\left\{\begin{array}[]{l}\bigl{(}\mathsf{vk},\,h,\,\{h_{i}\}_{i\in S_{R}},\\ \{\mathsf{ct}_{i}\}_{i\in S_{R}},\,\{\mathsf{ct}^{\prime}_{i}\}_{i\in S_{R}}\bigr{)},\\ \bigl{(}\mathsf{sk},\,\{x_{i}\}_{i\in S_{R}}\bigr{)}\end{array}\;\middle|\;\begin{array}[]{l}\mathsf{vk}=h^{\mathsf{sk}},\\[4.0pt] \forall i\in S_{R}:\;\mathsf{ct}_{i}=x_{i}+H(\mathsf{sk},i),\\[4.0pt] \forall i\in S_{R}:\;\mathsf{ct}^{\prime}_{i}=h_{i}^{\,\mathsf{sk}}\,g_{1}^{x_{i}}\end{array}\right\}.$
    6. 6.
      Output $\mathsf{(vk,sk,ct,\pi=\mathsf{(\pi_{Z},\pi_{R},ct^{\prime})}})$
  • $\mathsf{\mathsf{VECK^{\star}_{EL}}.Ver_{ct}(F_{[\ell]},C_{\phi},vk,ct,\pi)}\rightarrow 0/1$
    1. 1.
      Parse $\pi$ as $(\pi_{Z},\pi_{R},\mathsf{ct^{\prime}})$.
    2. 2.
      Compute $b_{1}\coloneqq\mathsf{VECK_{El}.Ver_{ct}(F_{S_{R}},C_{\phi},vk,ct^{\prime},\pi_{R})}$.
    3. 3.
      Compute $b_{2}\coloneqq\mathsf{Snark.Ver}(\pi_{Z})$.
    4. 4.
      Output $b_{1}\wedge b_{2}$.
  • $\mathsf{\mathsf{VECK^{\star}_{EL}}.Ver_{key}(vk,sk)}\rightarrow 0/1:$ For $\mathsf{sk}=s\in\mathbb{F}_{p}$, return 1 iff $\mathsf{vk}=h^{s}$.
  • $\mathsf{\mathsf{VECK^{\star}_{EL}}.Dec}(F_{[\ell]},sk,ct)\rightarrow\{\phi(i)\}_{i\in[\ell]}$
    1. 1.
      Compute the unmasked evaluations $\{\tilde{\phi}(i)\coloneqq\mathsf{ct}_{i}-H(\mathsf{sk},i)\}_{i\in[m]}$.
    2. 2.
      If $\mathsf{RS.Det}([m],\{\tilde{\phi}(i)\}_{i\in[m]})=0$ (no error detected), then output $\{\tilde{\phi}(i)\}_{i\in[\ell]}$. Otherwise, run the decoding algorithm and output
      $\mathsf{RS.Dec}\left([\ell],[m],\{\tilde{\phi}(i)\}_{i\in[m]}\right).$
Theorem 3.4
For the case $S=[\ell]$, the generation and verification times of $\mathsf{VECK^{\star}_{EL}}$ are $\mathcal{O}(\lambda)$.
Theorem 3.5
In the random-oracle and algebraic-group models, the protocol $\mathsf{\mathsf{VECK^{\star}_{EL}}}$ satisfies correctness, soundness, and computational zero-knowledge; hence it is a secure $\mathsf{VECK}$.
4 FDE via Payment Channels
Tas et al. [5] proposed a Bitcoin-based implementation of FDE that leverages adaptor signatures. However, their scheme utilizes pairing-friendly elliptic curves (e.g., BLS12-381), which are incompatible with the secp256k1 curve currently used by Bitcoin’s adaptor signatures. Consequently, direct integration of their proposed FDE protocols into Bitcoin faces significant practical barriers.
To resolve this incompatibility, we propose leveraging Bitcoin’s Lightning Network (LN) combined with zk-SNARKs. As an additional advantage, our approach naturally eliminates the need for costly on-chain transactions.
Specifically, the server (prover) generates a zk-SNARK proof $\pi_{t}$ attesting knowledge of a secret scalar $sk$ that satisfies the following relation:
$R\bigl{(}(h,vk,t),sk\bigr{)}=[vk=h^{sk}\wedge t=H(sk)],$
where $h$, $vk$, and $t$ are publicly known parameters. Although generating the zk-SNARK proof $\pi_{t}$ introduces computational overhead, this cost is constant with respect to the data size and can be precomputed.
Our resulting FDE protocol operates as follows:
  1. Step 1:
    The server sends to the client the ciphertext and its proof by calling $\mathsf{VECK^{+}_{El}.Enc}$, the value $t$, and the zk-SNARK proof $\pi_{t}$.
  2. Step 2:
    The client verifies all provided proofs, including $\pi_{t}$. Upon successful verification, the client initiates a payment via LN, using $t$ as the hash value in the HTLC.
Due to the atomicity provided by LN, the server obtains the payment iff it reveals the secret scalar $\mathsf{sk}$, enabling the client to decrypt the ciphertext. Crucially, this achieves atomic exchange without incurring additional costly blockchain transactions.
Fallback without LN.
If either party lacks LN connectivity, the same fairness guarantees can be enforced using exactly two Bitcoin transactions [^2] The original FDE protocol on Ethereum uses three on-chain transactions in the success path: registering the server’s public key, locking payment by the client, and the server revealing the decryption key [5]., structured as follows:
  1. 1.
    Funding Transaction: The client locks the payment amount in a P2WSH output with the script:
    $\mathtt{OP\_IF}$ $\mathtt{OP\_SHA256}~{}\langle t\rangle~{}\mathtt{OP\_EQUALVERIFY}$
    $\langle\mathsf{server\_pk}\rangle~{}\mathtt{OP\_CHECKSIG}$
    $\mathtt{OP\_ELSE}$ $\langle T_{\text{timeout}}\rangle~{}\mathtt{OP\_CHECKLOCKTIMEVERIFY}~{}\mathtt{OP\_DROP}$
    $\langle\mathsf{client\_pk}\rangle~{}\mathtt{OP\_CHECKSIG}$
    $\mathtt{OP\_ENDIF}$
    Here, $t=H(sk)$ is the hash commitment, and $T_{\text{timeout}}$ is a blockchain timestamp after which the client can reclaim the funds.
  2. 2.
    Spend or Refund Transaction: The above output is spent via one of two branches:
    • Success Path (before $T_{\text{timeout}}$): The server spends the output with witness data:
      $\langle sk\rangle,~{}\langle\sigma_{\text{server}}\rangle,~{}\mathtt{1}$
      The secret scalar $sk$ is thus revealed on-chain, allowing the client to decrypt the data.
    • Refund Path (after $T_{\text{timeout}}$): The client reclaims the output by providing:
      $\langle\sigma_{\text{client}}\rangle,~{}\mathtt{0}$
      No secret is revealed, and the client safely recovers their funds.
This fallback ensures fairness with minimal overhead when LN nodes are unavailable
Remark. An even more efficient fallback is to open a client-funded Lightning channel directly to the server. If the channel is closed immediately after a single purchase, this approach incurs exactly two on-chain transactions, matching the HTLC fallback described above. However, if the client expects additional purchases, subsequent file transfers incur virtually no extra latency or transaction fees, greatly reducing per-file amortized costs.
5 Conclusion
We presented a practical FDE instantiation that reduces proof generation and verification costs from linear to essentially constant in the security parameter $\lambda$. By leveraging Reed–Solomon coding with randomized sampling and introducing lightweight masked-helper encoding, we additionally reduced bandwidth overhead from at least 8×8$\times$8 × to approximately 2×2$\times$2 ×, while significantly accelerating client-side decoding. Formal security analysis confirms preservation of fairness, and experimental evaluation demonstrates substantial efficiency improvements. Overall, our results enhance the practicality and scalability of fair data exchange protocols.